How may revolutions per minute must the body make about vertical axis so that the cord make an angle of 45 with the vertical


1
Expert's answer
2022-03-22T15:48:17-0400

Answer

revolutions per second is given by

"=\frac{v}{2\pi R}\\=\frac{3.429}{2*3.14*1.2}\\=27.28rpm"